OpenStudy (anonymous):

If f(x) = x2 - x + 2, then find the value for f(-3). please show all your work i really need this its worth 10 percent of my grade

OpenStudy (raden):

it means ; f(-3) = (-3)^2 - (-3) + 2 = 9+3+2 = ...

OpenStudy (anonymous):

sorry im so bad but i dont understand:(

OpenStudy (raden):

given f(x) = x^2 - x + 2, and ur question is to find the value of f(-3). just put x=-3 into f(x), if u see everything of x u have to change by -3

OpenStudy (anonymous):

can u please just do it i dont understand and rea;ly need this i will fan u and give medal